# On Modular Invariants of Truncated Polynomial Rings in low ranks

### v1

@@ 3. A lower bound on the total dimension of the invariant subspaces
$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}=\delta_{2}(x^{q^{i}-1})+\text{essential monomials divisible by $Q_{2,0}^{\frac{q^{i+1}-1}{q-1}}$}.$
By repeating this process if necessary, we can write $f$ as a sum of essential monomials and Dickson polynomials of the form
$\delta_{2}(x^{q^{j}-1})Q_{2,1}^{a_{j}}Q_{2,0}^{b_{j}},\quad j\geq i.$
Thus it suffices to prove the lemma for $f=g\delta_{2}(x^{q^{i}-1})$, where $g\in{\mathcal{D}}_{2}$ and $10\leq i\leq m-$ 1.
To evaluate $\delta_{3}(g\delta_{2}(x^{q^{i}-1}))$, we first observe that by direct inspection, it can be simplified as a determinant:
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))=\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{2}^{q^{m}}&x_{3}^{q^{m}}\\ x_{1}^{q^{m}}g(x_{2},x_{3})&x_{2}^{q^{m}}g(x_{1},x_{3})&x_{3}^{q^{m}}g(x_{1},x_{2})\end{vmatrix}}{L_{3}(x_{1},x_{2},x_{3})}.$
Since
$L_{3}(x_{1},x_{2},x_{3})=$ $V_{3}(x_{1},x_{2},x_{3})L_{2}(x_{1},x_{2})$
this determinant can be rewritten in the form:
$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}h(x_{1},x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}h(x_{2},x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\bigg{]}.$
where
This quotient is a symmetric polynomial in $V_{2}^{q-1}(x_{2},x_{3})$ and $V_{2}^{q-1}(x_{2},x_{1})$ with coefficient in $\mathbb{F}_{q}[x_{2}^{q-1}]$. Moreover, if one summand of this quotient is divisible by $V_{2}(x_{2},x_{1})^{q-1}$, then the common denominator $L_{2}(x_{1},x_{2})$ will be canceled out, and the result is zero in ${\mathcal{Q}}_{m}$.
Thus if we write $g$ in terms of the upper triangular invariants, $g=\sum_{s,t}V_{1}^{s(q-1)}V_{2}^{t(q-1)}$, then the determinant above equals in ${\mathcal{Q}}_{m}$ to a sum over all such pair $(s,t)$, $t\geq 1$ of
$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}V_{2}^{(t-1)(q-1)}(x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}(x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\bigg{]},$
which is $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$. We can now use the computation in Lemma 7.2 to see that
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))\in\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{2}(Q_{1,0}^{a}),a\geq 0)\}.$
On the other hand, $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})$ is a genuine Dickon polynomial, hence $\delta_{3}(g\delta_{2}(x^{q^{i}-1}))$ is a $\operatorname{GL}_{3}$-invariant in ${\mathcal{Q}}_{m}$. This shows that
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))\in\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{3}(1)\}.$
The third, and final step is to replace the full Dickson algebra ${\mathcal{D}}_{3}$ by $\Delta^{m}_{3}$. We first need some preliminary calculations. In the following, we write $\delta_{3;a}(i_{1},i_{2})$ for $\delta_{3;a}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$ to simplify the notation. Again $\delta_{3}$ will stand for $\delta_{3;m}$.
@@
The following equalities hold:
@@
    $\delta_{3}(0,\frac{q^{i}-1}{q-1})=\delta_{3;m-i}^{q^{i}}(1)Q_{3,0}^{\frac{q^{i}-1}{q-1}}$ for all $i\geq 0$. In particular,
    $\delta_{3}(0,{\frac{q^{m-2}-1}{q-1}})=Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}\text{ and }\delta_{3}(0,\frac{q^{m-3}-1}{q-1})=Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}.$
@@
    $\delta_{3}(q^{m-3},\frac{q^{m-3}-1}{q-1})=Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$.
@@
    $\delta_{3}\bigg{(}\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1}\bigg{)}=\delta_{3;m-\lambda_{3}}^{q^{\lambda_{3}}}\bigg{(}\frac{q^{\lambda_{2}-\lambda_{3}}-1}{q-1},0\bigg{)}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}.$
@@
The proof of these statements are quite similar, and essentially makes use of the Laplace expansion. We will prove the last statement only. Apply the Laplace expansion along the last row of the determinant in the numerator for $\delta_{3}(\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1})$, we obtain
@@
The result follows easily after writing this expression as a $q^{\lambda_{3}}$ power. ∎
Our second technical result plays a similar role as Proposition 4.2 in the case of 3 variables.
@@
Suppose $20\leq\lambda\leq m-$ 2, then in ${\mathcal{Q}}_{m}$, $\delta_{3}(\frac{q^{\lambda}-1}{q-1},0)$ equals a Dickson polynomial in which the only non-essential summand which is not divisible by $Q_{3,0}$ is $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,1}^{\frac{q^{\lambda}-1}{q-1}}$.
@@
If $\lambda=0$, then $\delta_{3}(1)$ is a genuine Dickson polynomial in degree $q^{m}-q^{2}$. In this degree, there is only one Dickson monomial not in $\Delta^{m}_{3}$, namely $Q_{3,2}^{\frac{q^{m-2}-1}{q-1}}$. If we set $x_{2}=x_{3}=0$, then both $Q_{3,1}$ and $Q_{3,0}$ becomes zero, $\delta_{3}(1)$ becomes $x_{1}^{q^{m}-q^{2}}$. Thus in the expansion of $\delta_{3}(1)$ into a sum of Dickson monomials, $Q_{3,2}^{\frac{q^{m-2}-1}{q-1}}$ must occur with non-zero coefficient. The lemma is proved for $\lambda=0$.
We now assume that $\lambda\geq 1$. Note that $\delta_{3}(Q_{2,1})=\frac{[0,2,m]}{[0,1,2]}$ is a genuine Dickson polynomial, hence in ${\mathcal{Q}}_{m}$,
$\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})=Q_{3,2}^{\frac{q^{\lambda-1}-1}{q-1}}\delta_{3}(Q_{2,1})=Q_{3,2}^{\frac{q^{\lambda-1}-1}{q-1}}\frac{[0,2,m]}{[0,1,2]}.$
So $\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ equals in ${\mathcal{Q}}_{m}$ to a Dickson polynomial and we are going to use this representation.
Setting $x_{3}=0$, $\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ is reduced to
$\frac{[\lambda+1,m]}{[1,2]}=y^{q^{\lambda+1}}_{m-\lambda-1,0}Q_{2,0}^{\frac{q^{\lambda+1}-q}{q-1}}=Q_{2,1}^{\frac{q^{m-1}-q^{\lambda+1}}{q-1}}Q_{2,0}^{\frac{q^{\lambda+1}-q}{q-1}}+\text{others}.$
It follows that in this representation in ${\mathcal{Q}}_{m}$ of $\delta_{3;m}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ as a Dickson polynomial, $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,1}^{\frac{q^{\lambda-1}-1}{q-1}}$ must occur with nonzero coefficient. Moreover, any other monomials which involve $Q_{3,1}$ and $Q_{3,2}$ only must be divisible by $Q_{3,1}^{\frac{q^{\lambda}-1}{q-1}}$, and by degree reason, these monomials must be essential. ∎
@@
Suppose $m-3\leq\lambda_{2}\leq\lambda_{3}\leq 0$ and $\lambda\leq m-2$.
@@
    $\delta_{3}(\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1})=Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}+\text{other terms},$
    where the other terms are either essential or non-essential and divisible by $Q_{3,0}^{\frac{q^{\nu}-1}{q-1}}$, where $\nu>\lambda_{3}$.
@@
    $\delta_{3}(\frac{q^{m-2}-q^{\lambda}}{q-1},\frac{q^{\lambda}-1}{q-1})=Q_{3,1}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,0}^{\frac{q^{\lambda}-1}{q-1}}+\text{other terms},$
    where the other terms are either essential or non-essential and divisible by $Q_{3,0}^{\frac{q^{\nu}-1}{q-1}}$ where $\nu>\lambda$.
@@
This is just combining Lemma 7.4 and Proposition 7.5. ∎
@@
We have already seen from part (1) of Lemma 7.4 that $Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}$ belongs to $\Delta^{m}_{3}$. For the other 2 families, we will prove by downward induction on $\lambda_{3}$.
If $\lambda_{3}=m-3$, then $\lambda_{2}=\lambda_{3}=m-3$. The edge monomials $Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$ and $Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$ belong to the span of $\mathcal{B}_{m}(3)$ by the last 2 parts of Lemma 7.4. We now suppose that the results have been verified for all edge elements corresponding to $\lambda_{3}+1$ and above. Consider the edge elements corresponding to $\lambda_{3}$:
@@
Corollary 7.6 shows that these are in the image of $\delta_{3}(\Delta^{m}_{2})$ modulo those non-essential elements with strictly higher $\lambda_{3}$. By induction, these elements are already in the span of $\mathcal{B}_{m}(3)$. ∎
@@
Now we are ready to finish the proof. Corollary 7.7 shows that the subspace of ${\mathcal{Q}}_{m}^{G}(3)$ spanned by $\mathcal{B}_{m}(3)$ is closed under the action of the Dickson algebra ${\mathcal{D}}_{3}$. Indeed, we have already computed the action of ${\mathcal{D}}_{3}$ on $\delta_{s+1}^{3-s}(\Delta^{m}_{s})$ for $0\leq s\leq 2$ in 2.5, and the product of $Q_{3,i}$ with a monomial in $\Delta^{m}_{3}$ is clearly an edge monomial, which is in the span of $\mathcal{B}_{m}(3)$ by the above demonstration. The proof is finished.
8. A filtration of submodules
For each positive integer $n$, $0\leq k\leq n$, let $\mathcal{F}_{n,k}$ denote the $\mathbb{F}_{q}$-subspace of ${\mathcal{Q}}^{G}_{m}(n)$:
$\mathcal{F}_{n,k}=\operatorname{Span}(\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq\min(m,k)).$

### v2

@@ 3. A lower bound on the total dimension of the invariant subspaces
Repeating this process if necessary, we can write $f$ as a sum of essential monomials and Dickson polynomials of the form
$\delta_{2}(Q_{1,0}^{\frac{q^{j}-1}{q-1}})Q_{2,1}^{a_{j}}Q_{2,0}^{b_{j}},\quad j\geq i.$
Thus it suffices to consider $\delta_{3}(f)$ for $f=g\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})$, where $g\in{\mathcal{D}}_{2}$ and $10\leq i\leq m-$ 1.
We know from Corollary 7.3 that $\delta_{3}(g\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$ is a linear combination of polynomials in the following families:
  • $\delta_{2}^{2}(\Delta^{m}_{1})$,
  • $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{s(q-1)}$, $0\leq s\leq[m]_{q}$.
Since $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})$ is a genuine Dickson polynomial, $\delta_{3}(g\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$ is a $G_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$. We conclude that $\delta_{3}(g\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$ must belong to the space
$\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{3}(1)\},$
because any monomial of the form $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{i(q-1)}$, $0\leq i<[m]_{q}$, is not $G_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$. ∎
The third, and final step is to replace the full Dickson algebra ${\mathcal{D}}_{3}$ by $\Delta^{m}_{3}$. We first need some preliminary calculations. In the following, $\delta_{3}$ will again stand for $\delta_{3;m}$.
@@
The following equalities hold in $S$:
@@
    $\delta_{3}(Q_{2,0}^{\frac{q^{i}-1}{q-1}})=\big{(}\delta_{3;m-i}(1)\big{)}^{q^{i}}Q_{3,0}^{\frac{q^{i}-1}{q-1}}$ for all $i\geq 0$. In particular,
    $\delta_{3}(Q_{2,0}^{\frac{q^{m-2}-1}{q-1}})=Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}\text{\quad and \quad}\delta_{3}(Q_{2,0}^{\frac{q^{m-3}-1}{q-1}})=Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}.$
@@
    $\delta_{3}(Q_{2,1}^{q^{m-3}}Q_{2,0}^{\frac{q^{m-3}-1}{q-1}})=Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$.
@@
    $\delta_{3}\big{(}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{2,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}\big{)}=\bigg{(}\delta_{3;m-\lambda_{3}}\big{(}Q_{2,1}^{\frac{q^{\lambda_{2}-\lambda_{3}}-1}{q-1}}\big{)}\bigg{)}^{q^{\lambda_{3}}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}.$
@@
The proofs of these statements are quite similar and essentially make use of the Laplace expansion. We will prove the last statement only. Apply the Laplace expansion along the last row of the determinant in the numerator of $\delta_{3}\big{(}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{2,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}\big{)}$, we obtain
@@
After writing this expression as a $q^{\lambda_{3}}$ power, the result follows easily. ∎
Our second technical result is similar to Proposition 4.2 in the case of 3 variables.
@@
Let $m\geq 2$ be an integer.
  1. (i)
    The following equality holds in $S$:
    $\frac{[0,1,m-1]}{[0,1,2]}\;\delta_{3;m}(Q_{2,1})-\frac{[0,2,m-1]}{[0,1,2]}\;\delta_{3;m}(1)=Q_{3,1}^{\frac{q^{m-2}-1}{q-1}}+\text{other terms},$
    where the other terms are in the ideal $(Q_{3,0})$ of ${\mathcal{D}}_{3}$ generated by $Q_{3,0}$.
  2. (ii)
    For $20\leq\ell\leq m-$ 2, the following equality holds in $S$:
    $\frac{[0,1,\ell+1]}{[0,1,2]}\;\delta_{3;m}(Q_{2,1})-\frac{[0,2,\ell+1]}{[0,1,2]}\;\delta_{3;m}(1)=Q_{3,2}^{\frac{q^{m-2}-q^{\ell}}{q-1}}Q_{3,1}^{\frac{q^{\ell}-1}{q-1}}+\text{other terms},$
    where the other terms are in the ideal $(Q_{3,1}^{\frac{q^{\ell+1}-1}{q-1}},Q_{3,0})$ of ${\mathcal{D}}_{3}$.
@@
The key observation is that if two Dickson polynomials coincide after setting one of the variables, say $x_{3}$, to zero, then their difference must be a multiple of $Q_{3,0}$.
Note that in Dickson’s notation, we can write
$\delta_{3;m}(Q_{2,1})=\frac{[0,2,m]}{[0,1,2]},\quad\delta_{3;m}(1)=\frac{[0,1,m]}{[0,1,2]},$
so that all the terms in the above equations are genuine Dickson polynomials.
Proof of (i): When $x_{3}=0$, the equation under consideration becomes
$\frac{[1,m-1]}{[1,2]}\cdot\frac{[2,m]}{[1,2]}-\frac{[2,m-1]}{[1,2]}\cdot\frac{[1,m]}{[1,2]}=Q_{2,0}^{\frac{q^{m-1}-q}{q-1}}.$
It is now a simple computation that the two sides are the same.
Proof of (ii): Again, we consider the reduction of the left-hand side when setting $x_{3}=0$. The result, after simplification, is
$\delta_{2;m-\ell-1}(1)^{q^{\ell+1}}Q_{2,0}^{\frac{q^{\ell+1}-q}{q-1}}.$
Now setting $x_{2}=0$, $\delta_{2;m-\ell-1}(1)$ is reduced to $Q_{1,0}^{\frac{q^{m-\ell-1}-q}{q-1}}$, which implies that
$\delta_{2;m-\ell-1}(1)=Q_{2,1}^{\frac{q^{m-\ell-2}-1}{q-1}}+\text{other terms divisible by $Q_{2,0}$}.$
It follows that
$\delta_{2;m-\ell-1}(1)^{q^{\ell+1}}Q_{2,0}^{\frac{q^{\ell+1}-q}{q-1}}=Q_{2,1}^{\frac{q^{m-1}-q^{\ell+1}}{q-1}}Q_{2,0}^{\frac{q^{\ell+1}-q}{q-1}}+\text{otherterms divisible by $Q_{2,0}^{\frac{q^{\ell+2}-q}{q-1}}$.}$
Thus, we can conclude that the left-hand side equals
$Q_{3,2}^{\frac{q^{m-2}-q^{\ell}}{q-1}}Q_{3,1}^{\frac{q^{\ell}-1}{q-1}}+\text{other terms},$
where the other terms consist of Dickson monomials which are divisible by $Q_{3,1}^{\frac{q^{\ell+1}-1}{q-1}}$ or by $Q_{3,0}$. ∎
@@
Suppose $20\leq\ell\leq m-$ 2 and $m-3\geq\lambda_{2}\geq\lambda_{3}\geq 0$. The following equalities hold in $S$:
@@
    $\delta_{3;m-\ell-1}(1)^{q^{\ell}}\delta_{3;m}(Q_{2,1}^{q^{\ell}}Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})-\delta_{3;m-\ell-1}(Q_{2,1})^{q^{\ell}}\delta_{3;m}(Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})=Q_{3,1}^{\frac{q^{m-2}-q^{\ell}}{q-1}}Q_{3,0}^{\frac{q^{\ell}-1}{q-1}}+\text{other terms},$
    where the other terms are in the ideal $(Q_{3,0}^{\frac{q^{\ell+1}-1}{q-1}})$.
@@
    More generally,
    $\delta_{3;\lambda_{2}-\lambda_{3}+1}(1)^{q^{\lambda_{3}}}\delta_{3;m}(Q_{2,1}^{q^{\lambda_{3}}}Q_{2,0}^{\frac{q^{\lambda_{3}}-1}{q-1}})-\delta_{3;\lambda_{2}-\lambda_{3}+1}(Q_{2,1})^{q^{\lambda_{3}}}\delta_{3;m}(Q_{2,0}^{\frac{q^{\lambda_{3}}-1}{q-1}})=\\ Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}+\text{other terms},$
    where the other terms are in the ideal $(Q_{3,0}^{\frac{q^{\lambda_{3}+1}-1}{q-1}},Q_{3,1}^{\frac{q^{\lambda_{2}+1}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}-1}}{q-1}})$.
@@
This is just combining Lemma 7.5 and Proposition 7.6.
For the first part, note that
$\delta_{3;m}(Q_{2,1}^{q^{\ell}}Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})=\delta_{3;m}(Q_{2,1}^{\frac{q^{\ell+1}-q^{\ell}}{q-1}}Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})=\delta_{3;m-\ell}(Q_{2,1})^{q^{\ell}}Q_{3,0}^{\frac{q^{\ell}-1}{q-1}},$
and similarly
$\delta_{3;m}(Q_{2,0}^{\frac{q^{\ell}-1}{q-1}})=\delta_{3;m-\ell}(1)^{q^{\ell}}Q_{3,0}^{\frac{q^{\ell}-1}{q-1}}.$
The left-hand side of the required formula becomes
$\bigg{(}\delta_{3;m-\ell-1}(1)\delta_{3;m-\ell}(Q_{2,1})-\delta_{3;m-\ell-1}(Q_{2,1})\delta_{3;m-\ell}(1)\bigg{)}^{q^{\ell}}Q_{3,0}^{\frac{q^{\ell}-1}{q-1}}.$
We can then apply 7.6 part (i) to finish the proof.
The second part is proved similarly. ∎
@@
We have already seen from part (1) of Lemma 7.5 that the monomial $Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}$ belongs to $\delta_{3}(\Delta^{m}_{2})$. For the remaining $2$ families, we will show that they belong to the span of $\mathcal{B}_{m}(3)$ by downward induction on $\lambda_{3}$.
When $\lambda_{3}=m-3$, it follows that $\lambda_{2}=\lambda_{3}=m-3$. The edge monomials
$Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}\quad\text{and}\quad Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$
belong to the span of $\mathcal{B}_{m}(3)$ by parts (1) and (3) of Lemma 7.5. In addition, by Proposition 2.7, any Dickson monomial divisible by these monomials is also in the image of $\delta_{3}$, and thus belongs to the span of $\mathcal{B}_{m}(3)$.
Now suppose, as induction hypothesis, that all edge monomials with $Q_{3,0}$-exponent $\frac{q^{\lambda_{3}+1}-1}{q-1}$ as well as their multiples are already in the span of $\mathcal{B}_{m}(3)$. We consider the edge monomials corresponding to $\lambda_{3}$:
@@
Part (i) of Corollary 7.7 shows that $Q_{3,1}^{\frac{q^{m-2}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$ lies in the image of $\delta_{3}(\Delta^{m}_{2})$ modulo terms with strictly higher $\lambda_{3}$. By the induction hypothesis, these higher terms are already in the span of $\mathcal{B}_{m}(3)$.
For the second monomial, observe that if there exists a monomial with the same $Q_{3,0}$-exponent $\frac{q^{\lambda_{3}}-1}{q-1}$ but a higher exponent of $Q_{3,1}$, such that it becomes non-essential, then this monomial must be a multiple of
$Q_{3,1}^{\frac{q^{m-2}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}},$
which by the previous step, is already in the span of $\mathcal{B}_{m}(3)$. Alternatively, if the $Q_{3,0}$-exponent is strictly greater than $\frac{q^{\lambda_{3}}-1}{q-1}$, the induction hypothesis can be applied directly. ∎
@@
Now we are ready to finish the proof of the claim that $\mathcal{B}_{m}(3)$ is a generating set for ${\mathcal{Q}}_{m}(3)^{G}$. Corollary 7.8 shows that the subspace of ${\mathcal{Q}}_{m}(3)^{G}$ spanned by $\mathcal{B}_{m}(3)$ is closed under the action of the Dickson algebra ${\mathcal{D}}_{3}$. Indeed, we have already computed the action of ${\mathcal{D}}_{3}$ on $\delta_{s+1}^{3-s}(\Delta^{m}_{s})$ for $0\leq s\leq 2$ in 2.7, and the product of $Q_{3,i}$ with a monomial in $\Delta^{m}_{3}$ is clearly an edge monomial, which is in the span of $\mathcal{B}_{m}(3)$ by Corollary 7.8. The proof is finished.
8. A filtration of submodules over the Steenrod algebra
In this final section, we investigate the filtration $\mathcal{F}_{n,k}$ of ${\mathcal{Q}}_{m}(n)^{G}$ as defined in the Introduction 1.10. Recall that for each positive integer $n$, $0\leq k\leq\min(m,n)$, $\mathcal{F}_{n,k}$ is the $\mathbb{F}_{q}$-subspace of ${\mathcal{Q}}_{m}(n)^{G}$:
$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq\min(m,k)\}.$
The aim of this final section is to demonstrate Theorem 1.12. For $n\leq 3$, we have shown that this increasing filtration of subspaces is exhaustive. We begin with the following observation which allows more flexibility in working with $\mathcal{F}_{n,k}$:
Lemma 8.1.
Let $n\leq 3$. For each $0\leq k\leq\min(m,n)$, we have
$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in{\mathcal{D}}_{s},0\leq s\leq\min(m,k)\}.$
The point is that this filtration does not depend on the subspace $\Delta^{m}_{s}$.
Proof.
For $n=2$ this is essentially pointed out in Example 2.4 and the proof of Proposition 1.6. For $n=3$, this is a combination of Example 2.5 and the proof of Proposition 1.9. ∎
Corollary 8.2.
For $n\leq 3$, $\mathcal{F}_{n,*}$ is a filtration of $\mathcal{D}_{n}$-submodules of the invariant ring ${\mathcal{Q}}_{m}(n)^{G}$. Moreover, for $0\leq k<\min(m,n)$, $\mathcal{F}_{n,k}$ is annihilated by $Q_{n,0},\ldots,Q_{n,n-k-1}$.
Proof.
This is immediate from the above Lemma and Proposition 2.7. ∎